You are making a drink that requires 13 ounces of a 27-proof whiskey. Unfortunately, all you have on hand are 22-proof and 38-proof whiskey. How much of each type of whiskey do you need to mix to create 13 ounces of a 27-proof whiskey? The proof of an alcoholic beverage indicates the percentage of alcohol (ethanol) it contains. A 22-proof whiskey, for example, contains 44 percent alcohol by volume.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6

Answer by richwmiller(17219)

In the United States, proof is defined as double the percentage of alcohol contained in a solution at 60 degrees Fahrenheit or 15.6 degrees Celsius.
So it is the opposite of what you state.
200 proof is 100%
22 proof is 11%
38 proof is 19%
27 proof is 13.5%
We need a total of 13 oz. at 13.5% 27 proof
We want to know how many oz. at 11% 22 proof to mix with 19% 38 proof
a+b=13,
0.11*a+0.19*b=0.135*13
a=13-b
0.11*(13-b)+0.19*b=1.755
1.43-0.11b+0.19*b=1.755
0.08*b=0.325
b=4.0625
a=13-b
a=8.9375 oz. at 11% 22 proof
b=4.0625 oz. at 19% 38 proof
check
0.11*8.9375+0.19*4.0625=0.135*13
0.983125+0.771875=1.755
1.755=1.755
ok
codemx1